One electron has a charge of e=1.6x10^-19 coulombs. How many electrons must be transferred from an object for it to have a total charge of q=3.2x010^-3 coulombs?
Answer #1

Relation between number of electrons and charge is given by:

Q = n*e

Q = Charge on object = 3.2*10^-3 C

n = number of electrons = ?

e = charge on single electron = 1.6*10^-19 C

So,

Q = n*e

n = Q/e

n = 3.20*10^-3/(1.6*10^-19)

n = 2.0*10^16 electrons should be transferred

Please Upvote.